OpenStudy (anonymous):

cos 21 degrees = 9/y

zepdrix (zepdrix):

Solve for y? :o

zepdrix (zepdrix):

\[\Large \cos21^o=\frac{9}{y}\]So if we do a little algebra we can solve for y! :O Multiplying both sides by y gives us,\[\Large y \cos21^o=9\]We'll isolate the y term by dividing both sides by cos21,\[\Large y=\frac{9}{\cos 21^o}\]

zepdrix (zepdrix):

Understand how we did that? From here we would just want to use our old friend, Mr. Calculator, to get us the final value.